OpenStudy (anonymous):

there are 18 boys and 12 girls in math class. what is the ratio of girls to toatal student

OpenStudy (anonymous):

Add 18 + 12, which is the total amount of students, then you put the number of girls over that. Example:(x is the number of girls) x/18 + 12

OpenStudy (anonymous):

Total student = 18 + 12 = 30 no. of girls = 12 ratio of girls to total student = number of girls / total student ratio = 12 / 30 = 2/5

OpenStudy (anonymous):

@ashsek Don't give out the answer.

OpenStudy (anonymous):

@AnswerMyQuestions okay my fault

OpenStudy (anonymous):

Just don't do it again. :)

OpenStudy (anonymous):

okay sry

OpenStudy (anonymous):

@noterdame13 Add 12 + 18.